Part D The reactant concentration in a second-order reaction was 0.340 M after 280 s and 3.80x10-2 M after 850 s. What is the The reactant concentration In a second order reaction was 0.340 M After 280 seconds and 3.80×10^-2 M after 850 s. what is the rate constant for this reaction?
Answer #1

Second order react ron, | 1. Now, het 70.2401 = 5t CÁT, $ (280) + As = K (850) + 1 [A [3.80x1027 - ]substract ① from 63.90x1047 To. 3402 - @ 3(250) – 6C 220) 26.315 - 2.94 = O 23.375 (570) 23. 3.75 5.70 = 0.04100 = 4.1X10-2 Kthank you vmuch....